The gas, carbon dioxide, is found in dry air at sea level at a concentration of 3.30×10-2 percent by volume. What is the concentration of CO2 expressed in ppb? ppb
Answer #1

Answer :

3.30 * 10-2 percent by volume means 3.30 * 10-2 ml of CO​​​​​2 is present in 100 ml of sea water .

ppb = (volume of CO​​​​​2 / volume of solution ) * 109

= ( 3.3 * 10-2 / 100 ) * 109 = 3.3 * 105 ppb

Answer is 3.3 * 105 ppb concentration